If One sheepherder said to other sheepherder; give me ten of your sheep to have twice times of you. The other sheepherder answers him; give me ten of your sheep to have equals sheep. how many sheep has each one ?
Answer by stanbon(75887)

If One sheepherder said to other sheepherder; give me ten of your sheep to have twice times of you. The other sheepherder answers him; give me ten of your sheep to have equals sheep. how many sheep has each one ?
----
The two herders have A and B sheep respectively.
-----
Equations:
A+10 = 2(B-10)
A-10 = B+10
-------------------
Rearrange:
A - 2B = -30
A - B = 20
------------------
Subtract 1st from 2nd and solve for "B"
B = 50 sheep
----
Since A = B+20, A = 70 sheep
===============================
Cheers,
Stan H.